OpenStudy (anonymous):

HELP GIVING MEDALS N BECOMING A FAN!!!!!! 1. A capacitor has a capacitance of 3 pF. What potential difference would be required to store a charge of 21 pC? 3 V 21 V 7 V 0.14 V 2. A capacitor of capacitance 4 μF is connected to a 1.5 V battery. How much electrical potential energy is stored? 9.0 x 10^-6 J 2.7 x 10^-6 J 6.0 x 10^-6 J 4.5 x 10^-6 J

OpenStudy (anonymous):

@IrishBoy123 @DanJS @SolomonZelman @Nnesha

OpenStudy (irishboy123):

for first question, we know capacitance C as: \(\huge C = \frac{Q}{V}\). for second question, this is the energy equation in its 3 forms \(\huge E = \frac{Q^2}{2\ C} = \frac{C\ V^2}{2} = \frac{Q \ V}{2}\) you should have a play, choose the right one for the situation, and see where you go.